OpenStudy (anonymous):

The manager of a large apartment complex knows from experience that 110 units will be occupied if the rent is 324 dollars per month. A market survey suggests that, on the average, one additional unit will remain vacant for each 9 dollar increase in rent. Similarly, one additional unit will be occupied for each 9 dollar decrease in rent. What rent should the manager charge to maximize revenue? This is an optimization problem and I have a really hard time doing these so if someone could help it would be great! Thank you! :)

OpenStudy (anonymous):

\[R = revenue\] \[R = (110-x)(324+9x) \ge 0\ \\x \epsilon (0, 110)\] \[R' = -18(x-37)\] Now test: When R'=0, x=37 When R'=DNE, it's impossible Now using second derivative test to check if x=37 is a local max or local min: \[R''=-18\] This is a negative number, which means it's a local max. Hence, rent is \[342+9x=342+9\times18=504$\]

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

Try 657

OpenStudy (anonymous):

yup works

OpenStudy (anonymous):

Thank you :) can you look at my other question I posted and I'll post my last one right now if you don't mind helping me with those :)

OpenStudy (anonymous):

Alright